In an open economy,

(a) I = S + EX

(b) I = S + CA

(c) I = S - EX

(d) I = S + IM - EX

(e) I = S + EX - IM
Answer #1

Answer

In an open economy, Economy is in equilibrium when Y = AE where AE = aggregate expenditure = C + I + G + EX - IM, Y = GDP, C = consumption, I = investment, G = government spending, EX = exports and IM = Imports.

Hence we have Y = C + I + G + EX - IM => I = Y - C - G - EX + IM --------------------(2)

National saving(S) = Private saving + Public saving, where Private saving = Y - C - T where T = taxes and Public saving = T - G

=> S = (Y - C - T) + (T - G) = Y - C - G -----------------------(2)

From (1) and (2) we get :

I = Y - C - G - EX + IM = S - EX + IM

=> I = S + IM - EX

Hence, the correct answer is (d) I = S + IM - EX